Required information Use the following information for the Problems below. [The following information applies to the questionRequired: 1. Compute the direct materials cost variance, including its price and quantity variances. Actual Cost Standard Cos2. Compute the direct labor cost variance, including its rate and efficiency variances. Actual Cost Standard Cost3. Compute the overhead controllable and volume variances. Controllable Variance Actual overhead Budgeted overhead Controllab
Answer #1

1. Direct Material Cost Variance

Actual Quantity = 1,466,000

Standard Quantity = 1,863,000

Actual Price = $ 7.4

Standard Price = $ 4.4

Material Price Variance =Actual Quantity x Actual Price - Actual Quantity x Standard Price

Material Price Variance = (1466000 * 7.4) - (1466000 * 4.4)

= 10,848,400-6,450,400

=4,398,000(Adverse)

Material Quantity Variance = Actual Quantiy x Standard Price - Standard Quantity x Standard Price

= (1466000 * 4.4) -( 1863000 * 4.4)

=6,450,400 - 8,197,200

= (1,746,800) (Favourable)

Material Cost Variance = Material Price Variance + Material Quantity Variance

= 4,398,000 (A) + (1,746,800) (F)

= 2,651,200 (A)

2. Direct Labour Cost Variance

Standard Direct Labor Hour = 372600

Actual Direct Labor Hour = 291000

Standard Rate/Hour = $ 14/Hr

Actual Rate/Hour = $12.6/Hr

Labor Rate Variance = Actual Labor Hours * Actual Rate - Actual Labor Hours * Standard Rate

= (291000 * 12.6) - (291000 * 14)

= 3,666,600 - 4,074,000

= -407,400 (Favourable)

Labor Efficiency Variance = Actual Hours x Standard Rate - Standard Hours x Standard Rate

= 291000* 14 - 372600* 14

= 4,074,000- 5,216,400

= -1,142,400 (Favourable)

Direct Labor Variance = Labor Rate Variance + Labor Efficiency Variance

= (- 407,400) (f) + (-1,142,400) (f)

= -1,549,800 (Favourable)

3. Overhead Controllable Variance = Actual Factory Overhead-Budgeted Factory overhead

Actual Factory Overhead = 3014600 +3272200 = $ 6,286,800

Budgeted Allowance Based on standard Hours Allowed = $ 7452000

Overhead Controllable variance=6,286,800-7452000

= −1,165,200 Favorable

Overhead Volume Variance = Actual Fixed Cost Absorbed - Budgeted Fixed Cost

= 3014600- 3974400

= - $959,800(Favourable)

5 Direct Material Cost Variance X AQ 1,466,000 SP $4.40 SQ 1863000 Standard cost X I | 8.197,200 SP $4.40 | 10 4.398,000(Adve